Rosa works due north of home. Her husband Ray works due east. They leave for work at the same time. By the time Rosa is 5 miles from home, the distance between them is one mile more more than Ray's distance from home. How far from home is Rosa?

Answer by fractalier(6550)

You can construct a right triangle...
Rosa's leg is 5 miles.
Call Ray's leg, x.
The hypotenuse is then x + 1.
From the Pythagorean Theorem, we have
x^2 + 5^2 = (x + 1)^2
x^2 + 25 = x^2 + 2x + 1
2x = 24
x = 12
Rosa is still 5 miles from home.
Ray is 12 miles from home.
Check the question.